# Emergent SU(3) topological system in a trimer SSH model

### v1

@@ III Topological phases and bulk-boundary correspondence
where $H(k)$ is the Bloch Hamiltonian. The winding number can only be quantized and is $\mathbb{Z}$-valued for systems respecting chiral symmetries.
@@
For inversion-symmetric topological insulators with or without chiral symmetry, the Zak phase has been used to distinguish topologically trivial ($\phi_{1}=0~{}\rm mod~{}2\pi)$ and non-trivial phases ($\phi_{1}=\pi~{}\rm mod~{}2\pi$). However, there is an ambiguity in establishing the bulk-boundary correspondence. In a recent study [47], a generalization of the Zak phase namely the inter-cellular Zak phase has been proposed which establishes the bulk-boundary correspondence by correctly counting the number of boundary modes below the Fermi level. The inter-cellular Zak phase is the sum of sub-lattice Zak phases [48] which is defined as follows
@@
where $P_{\alpha}$ denotes the projection operator on the sub-lattice $\alpha$. In the next two sub-sections, we will discuss bulk-boundary correspondence and two different topological invariants namely the Zak phase and sub-lattice winding numbers.
Figure 4: Bulk-topological invariant: (a), (b), and (c) describe the Zak phases ($\phi_{1},~{}\phi_{2},~{}\text{and}~{}\phi_{3}$ in units of $\pi$) for $\lambda=1,~{}\lambda=2,$ $\lambda=3$ bands, respectively, in the $(w,w_{1})$-plane for $u=v=1$. (d) The sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$-plane for $u=v=1$.
Figure 5: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping amplitude $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=0$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.52,0.0)$. (d) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.52,0.0)$.
Figure 6: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping term $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=2$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.72,2.0)$. (d) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.72,2.0)$.
III.1 Bulk-boundary correspondence in trimer SSH-model
Recently, the bulk-boundary correspondence in the trimer SSH model [42, 43] without NNN hopping terms has been studied. The trimer SSH model lacks full chiral symmetry but has generalized chiral, isolated, and point chiral symmetries. Therefore, the chiral winding number cannot characterize the topological phases in this system. However, the Zak phase successfully establishes the bulk-boundary condition for the system parameters respecting the inversion symmetry. Moreover, for the trimer SSH model in the absence of NNN hopping term with broken inversion symmetry, another generalization of the Zak phase namely normalized sub-lattice Zak phase has been proposed which successfully establishes the bulk-boundary condition [43].
In this section, we discuss the bulk-boundary condition in the inversion-symmetric trimer SSH lattice with NNN hopping terms ($u_{1},~{}v_{1},~{}w_{1}$) (Fig. 1). In particular, the inversion symmetry enforces the hopping to satisfy the following criteria,
$\Delta=0,u=v,~{}\text{and}~{}u_{1}=v_{1}.$ (21)
For simplicity, we will consider $u=v=1$ which sets the unit of energy and $u_{1}=v_{1}=0$ in this section. The results for different sets of $u_{1}=v_{1}\neq 0$ is given in the Appendix C.
$w_{1}=0$ case: In the absence of NNN coupling ($w_{1}=0$) and for $u=v=w$, the lower band touches the middle one at $\Gamma$-point ($k=0$) and the the middle band touches the upper one at $X$-point ($k=\pi$) simultaneously, signaling a topological phase transition (See Fig 2).
$w_{1}\neq 0$ case: For finite NNN hopping ($w_{1}\neq 0$), the degeneracy of the band touching points at $\Gamma,~{}X$-points is removed. In this case, the bulk band structure evolves according to Fig. 3 as we tune the inter-cell hopping $w$ for the fixed value of NNN hopping $w_{1}=2$. At $w=u=v$ and $w<w_{1}$, the band gap between the lower and middle bands closes at the $X$-point (See Fig.  3 (b)). Similarly, at $w>u=v$ and $w=w_{1}$, the band gap between the middle and the upper bands closes at the $X$-point (See Fig.  3 (d).
We numerically, compute the Zak phase defined in Eq. 19 for the individual bands. Figures 4 (a), (b), (c), and (d) show the variation of the Zak phases $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, and $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the ($w,w_{1}$)-plane, respectively. In the absence of NNN hopping ($w_{1}=0$), our results are in full agreement with that of Ref. [42, 43]. It can be seen that any two of the three Zak phases change in a quantized manner whenever the bulk band gap closes, which signals a topological phase transition. Accordingly, in Table 1, we label all three different topological phase transitions (topological phase transitions) characterized by the locations of the gap-closing points in the BZ while tuning $w$ and $w_{1}$.
@@ III.1 Bulk-boundary correspondence in trimer SSH-model
TPT-I $w>u=v,w_{1}=0$ $u_{1}=v_{1}=0$ $\Delta_{21}$($\Delta_{32}$) closes at $\Gamma$ ($X$)-point $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ $\nu^{A}=0\leftrightarrow 2$
TPT-II $w>u=v,w<w_{1}$ $u_{1}=v_{1}\neq 0$ $\Delta_{21}$ closes only at $X$-point $\phi_{1}+\phi_{3}=0\leftrightarrow~{}\pi$ $\nu^{A}=0\leftrightarrow 1$
TPT-III $w>u=v,w>w_{1}$ $u_{1}=v_{1}\neq 0$ $\Delta_{32}$ closes only at $X$-point $\phi_{1}+\phi_{3}=\pi\leftrightarrow 2\pi$ $\nu^{A}=1\leftrightarrow 2$
@@
Table 1: Classification of topological phase transition (TPT): Three topological phase transitions for various system parameters based on the band gaps $\Delta_{21}=E_{2}-E_{1}$ and $\Delta_{32}=E_{3}-E_{2}$ closing points in the BZ along with the topological invariant quantities such as Zak phase and sub-lattice winding number
For $w>u=v$ and $w_{1}=0$, the Zak phases $\phi_{1},\phi_{3}$ are quantized to $\pi$ while $\phi_{2}$ vanishes. In the open boundary condition, there exists at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system [see Fig. 5 (a) ]. These results establish the bulk-boundary correspondence. In this case the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are the same and are determined by the bulk energy gaps $\Delta_{21}$ and $\Delta_{32}$. Figure 5 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figures  5 (c) and (d) show the real space profile of the wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively.
Figures 6 (a) and (b) show the variation of the band structure in the open boundary condition and $\phi_{1}+\phi_{3}$ with respect to $w$, respectively. Figures 6 (c) and (d) show the real space profile of wave function amplitudes for a finite open chain corresponding to the boundary modes with energies $E_{\text{b}}=+1$ and $E_{\text{b}}=-1$, respectively. Accordingly, in the open boundary condition, there exist at each energy $E_{\text{b}}=\pm 1$, a single boundary mode localized at each end of the system. In this case the localization lengths for the boundary modes at energies $E_{\text{b}}=\pm 1$ are different and are determined by the bulk energy gap $\Delta_{21}$ ($\Delta_{32}$) for negative (positive) energy boundary modes.
Based on the above analysis we observe that the addition of Zak phases $\phi_{1}+\phi_{3}$ characterizes the topological phase transitions and number of boundary modes at each end of the system. A similar analysis holds for $u_{1}=v_{1}\neq 0$. However, the phase diagram looks more complex, which is given in Appendix C (See Fig. C.7).
III.2 Sub-lattice winding numbers as bulk-topological invariants
Here we propose a different bulk topological invariant namely the sub-lattice winding number, similar to the sub-lattice Zak phases defined in Eq. 20, to discuss the bulk-boundary correspondence in this system. We construct the doubled Hermitian Hamiltonian from the chiral asymmetric Bloch Hamiltonian $H(k)$ in the following manner [49, 50]
$\tilde{H}(k)=\begin{pmatrix}0&\mathcal{H}(k)-E_{\rm b}\\ \mathcal{H}^{\dagger}(k)-E_{\rm b}&0\end{pmatrix},$ (24)
which by construction is chiral symmetric with respect to the following chiral symmetry operator
$\Sigma_{z}=\begin{pmatrix}I_{3}&0\\ 0&-I_{3}\end{pmatrix},$ (27)
where $I_{3}$ denotes the $3\times 3$ identity matrix. We choose the reference energy $E_{\rm b}$ same as the gap closing energy. The gaped nature of $\tilde{H}_{E_{\rm b}}(k)$ is defined via $\det(\tilde{H}_{E_{\rm b}}(k))\neq 0$ which is true for $\det(\mathcal{H}(k)-E_{\rm b})\neq 0$. For non-Hermitian systems, the symmetry class chiral symmetry is called class AIII, and the topological classification of insulators is given by the bulk winding numbers $\nu\in\mathbb{Z}$ which is defined as follows
$\nu_{E_{\rm b}}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[(\mathcal{H}(k)-E_{\rm b})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm b})].$ (28)
It is straightforward to realize that the above topological invariant vanishes for the present case. However, we can define the sub-lattice winding numbers $\nu_{E_{\rm b}}^{\alpha}$ in the following manner,
$\nu_{E_{\rm b}}^{\alpha}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[P_{\alpha}(\mathcal{H}(k)-E_{\rm b})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm b})P_{\alpha}],$ (29)
where $P_{\alpha}=|\alpha\rangle\langle\alpha|$ denote the projection operators for different sub-lattice sectors $\alpha=A,B,C$. Here the sub-lattice eigenstate can be expressed as $|\alpha\rangle=(\delta_{\alpha,A},\delta_{\alpha,B},\delta_{\alpha,C})^{\mathcal{T}}$ with ${\mathcal{T}}$ being the transpose operator. Although, the total winding number ($\nu_{E_{\rm b}}=\sum_{\alpha}\nu_{E_{\rm b}}^{\alpha}=0)$ vanishes for each reference energy. The sub-lattice winding numbers ($\nu_{E_{\rm b}}^{\alpha}$) can be quantized and may characterize the number of boundary modes in the system. This is indeed the case for the present system. We construct a bulk topological invariant $\nu^{\rm A}$ which is defined as follows
$\nu^{\rm A}=\sum_{E_{\rm b}}\nu_{E_{\rm b}}^{\rm A}.$ (30)
The non-trivial value of the above topological invariant $\nu^{\rm A}$ counts the number of boundary modes localized at each end of the system and establishes the bulk-boundary correspondence.

### v2

@@ III Topological invariants
where $H(k)$ is the Bloch Hamiltonian. The winding number can only be quantized and is $\mathbb{Z}$-valued for systems respecting chiral symmetries. It is to be noted that the bulk winding number for non-chiral system is always zero.
@@
Figure 4: Bulk-topological invariant: (a), (b), and (c) describe the Zak phases ($\phi_{1},~{}\phi_{2},~{}\text{and}~{}\phi_{3}$ in units of $\pi$) for $\lambda=1,~{}\lambda=2,$ $\lambda=3$ bands, respectively, in the $(w,w_{1})$-plane for $u=v=1$. (d) The sum of the Zak phases $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the $(w,w_{1})$-plane for $u=v=1$. Different lines mark the topological phase transitions. Horizontal dotted lines for $w_{1}=0$ and $w_{1}=2.0$ show the parameters of Fig. 6 and Fig. 7, respectively.
Figure 5: Bulk-topological invariant: the A sub-lattice winding number. (a) and (b) show the behavior of sub-lattice winding number ($\nu_{\rm A}$) in the $(w,w_{1})$-plane for the inversion symmetric ($u=v=1,u_{1}=v_{1}=0$) and inversion asymmetric ($u=1,v=2,u_{1}=v_{1}=0$) chain, respectively. In (a) and (b), different lines and numbers mark the topological phase transition lines and the number of topological edge states that coincide with the sublattice winding number, respectively. In panel (a), horizontal dotted lines for $w_{1}=0$ and $w_{1}=2.0$ show the parameters of Fig. 6 and Fig. 7, respectively. In panel (b), horizontal dotted lines for $w_{1}=0$ and $w_{1}=0.2$ show the parameters of Fig. 8 and Fig. 9, respectively.
For inversion-symmetric topological insulators with or without chiral symmetry, the Zak phase has been used to distinguish topologically trivial ($\phi_{1}=0~{}\rm mod~{}2\pi)$ and non-trivial phases ($\phi_{1}=\pi~{}\rm mod~{}2\pi$). However, there is an ambiguity in establishing the bulk-boundary correspondence, as the Zak phase itself depends on the real space origin and choice of unit cell required to define the Bloch Hamiltonian. In a recent study [53], a generalization of the Zak phase, namely the inter-cellular Zak phase, has been proposed, which establishes the bulk-boundary correspondence by correctly counting the number of boundary modes below the Fermi level. The inter-cellular Zak phase is the sum of sub-lattice Zak phases [54] which is defined as follows
@@
where $P_{\alpha}=|\alpha\rangle\langle\alpha|$ denotes the projection operator on the sub-lattice $\alpha$ with $|\alpha\rangle$ being the sub-lattice eigenstate.
Here we propose a different bulk topological invariant namely the sub-lattice winding number, similar to the sub-lattice Zak phases defined in Eq. 18, to discuss the bulk-boundary correspondence in this system. We emphasize here that there are two advantages of defining sub-lattice winding numbers. Firstly, the sub-lattice winding number gives a more straightforward understanding of the topological phases, similar to the winding number for the dimer SSH model [2]. Secondly, we show that the sub-lattice winding number can still be quantized for the inversion asymmetric chain and successfully establish the bulk-boundary correspondence in the system.
Our system has no non-trivial topological phases as it belongs to class AI of the 10-fold AZ classification of the Hermitian system [9, 10]. The point gap-topological phases in non-Hermitian systems also have no non-trivial topological phases according to the 10-fold AZ classification of the Hermitian system. However, the topology of the point-gap non-Hermitian systems can be understood from the point of view of the topology of the doubled Hermitian Hamiltonian [see Refs. [55, 56] for more details on 38-fold AZ classification of the non-Hermitian system]. For point-gap topological phases, the winding number characterizes the point-gap topology for a given reference eigenvalue $E_{\rm r}$ and Bloch Hamiltonian $\mathcal{H}(k)$, as given by [55, 56]
$\nu_{E_{\rm r}}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[(\mathcal{H}(k)-E_{\rm r})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm r})]$ (19)
The non-trivial values of $\nu_{E_{\rm r}}\in\mathbb{Z}$ characterize the point gap topology of the systems and correspond to the edge states (skin modes). This type of non-Hermitian topology has no Hermitian analog and can not be understood from the 10-fold AZ classification of Hermitian systems. However, it can be understood by defining a doubled Hermitian Hamiltonian
$\tilde{H}_{E_{\rm r}}(k)=\begin{pmatrix}0&\mathcal{H}(k)-E_{\rm r}\\ \mathcal{H}^{\dagger}(k)-E_{\rm r}&0\end{pmatrix},$ (20)
which by construction has chiral symmetry for the following chiral symmetry operator
$\Sigma_{z}=\begin{pmatrix}I_{3}&0\\ 0&-I_{3}\end{pmatrix},$ (21)
where $I_{3}$ denotes the $3\times 3$ identity matrix and belongs to the Class AIII of the 10-fold AZ classification. It is straightforward to see that the winding number $\nu_{E_{\rm r}}$ (winding of the off-diagonal element of ~$\tilde{\mathcal{H}}_{E_{\rm r}}(k)$) also characterizes the topology of the ~$\tilde{\mathcal{H}}_{E_{\rm r}}(k)$. Therefore, the skin modes with reference eigenvalue $E_{\rm r}$ correspond to a pair of gapless topological edge modes under the open boundary condition.
Following the above arguments, we construct the doubled Hermitian Hamiltonian $\tilde{H}_{E_{\rm b}}(k)$ from the chiral asymmetric Bloch Hamiltonian $H(k)$ via replacing the reference energy $E_{\rm r}$ to the gap-closing energy $E_{\rm b}$. The gaped nature of $\tilde{H}_{E_{\rm b}}(k)$ is defined via $\det(\tilde{H}_{E_{\rm b}}(k))\neq 0$ which is true for $\det(\mathcal{H}(k)-E_{\rm b})\neq 0$. As our system is Hermitian, it is straightforward to realize that the winding number $\nu_{E_{\rm b}}$ vanishes, as expected. However, we can define the sub-lattice winding numbers $\nu_{E_{\rm b}}^{\alpha}$ in the following manner,
$\nu_{E_{\rm b}}^{\alpha}$ $=\oint_{\rm BZ}\frac{dk}{2\pi i}{\rm Tr}[P_{\alpha}(\mathcal{H}(k)-E_{\rm b})^{-1}\partial_{k}(\mathcal{H}(k)-E_{\rm b})P_{\alpha}].$ (22)
The sub-lattice eigenstate for the trimer SSH model can be expressed as $|\alpha\rangle=(\delta_{\alpha,A},\delta_{\alpha,B},\delta_{\alpha,C})^{\mathcal{T}}$ with ${\mathcal{T}}$ being the transpose operator. Although the total winding number ($\nu_{E_{\rm b}}=\sum_{\alpha}\nu_{E_{\rm b}}^{\alpha}=0)$ vanishes for each reference energy. The sub-lattice winding numbers ($\nu_{E_{\rm b}}^{\alpha}$) can be quantized and may characterize the number of boundary modes in the system. This is indeed the case for the present system. We construct a bulk topological invariant $\nu^{\rm A}$ which is defined as follows
$\nu^{\rm A}=\sum_{E_{\rm b}}\nu_{E_{\rm b}}^{\rm A}.$ (23)
The non-trivial value of the above topological invariant $\nu^{\rm A}$ counts the number of boundary modes localized at each end of the system and establishes the bulk-boundary correspondence. The physical implication of the zero winding number for the B sub-lattice and equal and opposite winding numbers for the A and C sub-lattices is related to the edge states profile. In the topologically non-trivial phases, the winding number for the A sub-lattice takes finite quantized values that are equal to and opposite to that of the C sub-lattice. The physical consequence is that the edge states have zero wave function amplitudes at the A and C sub-lattice sites for the right and left boundaries of the system, respectively. However, the winding number of the B sub-lattice vanishes, implying the non-vanishing of the wave function amplitudes at both the boundaries of the systems [See Figs. 6-9]. This is also the case for the dimer SSH model, where the finite winding number for both sub-lattices A and B corresponds to the vanishing of the wave function amplitudes for the boundary modes at the A and B sub-lattice sites for the right and left boundaries of the system, respectively. In the Appendix E, we discuss a detailed analysis of the real space profile of the edge states profile.
We note here that the gap closing energy $E_{\rm b}$ does not depend on the size of the system, rather it depends on the system parameters in the Bloch Hamiltonian, e.g. $E_{\rm b}=\pm u$, for $u=v$. Therefore, the sublattice winding number $\nu_{\rm A}$ [Eq. (23)] can be used to establish the bulk-boundary correspondence.
@@
TPT-I $w=|u|,w_{1}=u_{1}=v_{1}=0$ Both $\Delta_{21}$, $\Delta_{32}$ close. $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ $\nu^{A}=0\leftrightarrow 2$
TPT-II $w=|w_{1}-u|,w_{1},u_{1}=v_{1}\neq 0$ Only $\Delta_{21}$ closes. $\phi_{1}+\phi_{3}=0\leftrightarrow~{}\pi$ $\nu^{A}=0\leftrightarrow 1$
TPT-III $w=|w_{1}+u|,w_{1},u_{1}=v_{1}\neq 0$ Only $\Delta_{32}$ closes. $\phi_{1}+\phi_{3}=\pi\leftrightarrow 2\pi$ $\nu^{A}=1\leftrightarrow 2$
@@
Table 1: Classification of topological phase transition (TPT) for inversion symmetric system ($u=v$): Three topological phase transitions for various system parameters based on the band gaps $\Delta_{21}=E_{2}-E_{1}$ and $\Delta_{32}=E_{3}-E_{2}$ closing points along with the topological invariant quantities such as Zak phase and sub-lattice winding number [see Figs. 4, 5, and C.10].
Figure 6: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping amplitude $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=0$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.52,0.0)$. (d) The real space profile of the probability amplitudes (Inset: wave function amplitudes) corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.52,0.0)$.
Figure 7: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping term $w$ in the open boundary condition. The horizontal dashed lines are the edge modes with energy $E_{\text{b}}=\pm 1$. (b) Bulk topological invariant (sum of the Zak phases for the two bands $\phi_{1}+\phi_{3}$ (in units of $\pi$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=v=1$, and $w_{1}=2$. (c) The real space profile of the probability amplitude (Inset: wave function amplitudes) corresponding to the boundary mode at $E_{\text{b}}=+1$ for $(w,w_{1})=(3.72,2.0)$. (d) The real space profile of the wave function amplitudes corresponding to the boundary mode at $E_{\text{b}}=-1$ for $(w,w_{1})=(3.72,2.0)$.
Figure 8: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping term $w$ in the open boundary condition. The horizontal highlighted lines are the edge modes with energies $E_{\text{b}}=\pm u$ and $E_{\text{b}}=\pm v$, which appear at mirror symmetric points $w=\pm u$ and $w=\pm v$, respectively. (b) Bulk topological invariant (sub-lattice winding number $\nu_{\rm A}$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=1,v=2$, and $w_{1}=u_{1}=v_{1}=0$. (c) The real space profile of the probability amplitudes (Inset: wave function amplitudes) corresponding to the boundary mode at $E_{\text{b}}=\pm u$ for $(w,w_{1})=(3.52,0.0)$. (d) The real space profile of the probability amplitudes (Inset: wave function amplitudes) corresponding to the boundary mode at $E_{\text{b}}=\pm v$ for $(w,w_{1})=(3.52,0.0)$.
Figure 9: Bulk-boundary correspondence: (a) Energy band structure vs inter-cell hopping term $w$ in the open boundary condition. The horizontal highlighted lines are the edge modes with energies $E_{\text{b}}=\pm u$ and $E_{\text{b}}=\pm v$ which appear at $w=\pm(w_{1}\pm u)$ and $w=\pm(w_{1}\pm v)$, respectively. (b) Bulk topological invariant (sub-lattice winding number $\nu_{\rm A}$) characterizes the appearance of boundary modes and hence the topological phase transition as one tunes the inter-cell hopping term $w$ for $u=1,v=2$, and $w_{1}=0.2,u_{1}=v_{1}=0$. (c) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=\pm u$ for $(w,w_{1})=(3.52,0.2)$. (d) The real space profile of the wave function amplitude corresponding to the boundary mode at $E_{\text{b}}=\pm v$ for $(w,w_{1})=(3.52,0.2)$.
IV Topological phases and bulk-boundary correspondence in the absence of NNN hopping
Recently, the bulk-boundary correspondence in the trimer SSH model [47, 48] without NNN hopping terms has been studied. The trimer SSH model lacks full chiral symmetry but has generalized chiral, isolated, and point chiral symmetries. Therefore, the chiral winding number cannot characterize the topological phases in this system. However, the Zak phase successfully establishes the bulk-boundary correspondence for the system parameters respecting the inversion symmetry. Moreover, for the trimer SSH model in the absence of NNN hopping term with broken inversion symmetry, another generalization of the Zak phase, namely normalized sub-lattice Zak phase has been proposed, which successfully establishes the bulk-boundary correspondence [48]. In this section, we discuss the bulk-boundary correspondence in the absence of NNN hopping terms ($u_{1},~{}v_{1},~{}w_{1}$) (Fig. 1).
IV.1 Bulk case
$w_{1}=0$ case: In the absence of NNN coupling ($w_{1}=0$) and for $u=v=w$, the lower band touches the middle one at $\Gamma$-point ($k=0$) and the the middle band touches the upper one at $X$-point ($k=\pi$) simultaneously signaling a topological phase transition (See Fig 2).
We numerically compute the Zak phase defined in Eq. 17 for the individual bands. Figures 4 (a), (b), (c), and (d) show the variation of the Zak phases $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, and $\phi_{1}+\phi_{3}$ (in units of $\pi$) in the ($w,w_{1}$)-plane, respectively. In the absence of NNN hopping ($w_{1}=0$), our results are in full agreement with that of Ref. [47, 48]. It can be seen that any two of the three Zak phases change in a quantized manner whenever the bulk band gap closes, which signals a topological phase transition. Accordingly, in Table 1, we label all three different topological phase transitions characterized by the locations of the gap-closing points in the BZ while tuning $w$ and $w_{1}$.